A 10% (NH4)2SO4 standard is needed. How much solute would be needed to make 50mL?
1
Expert's answer
2019-11-05T07:39:36-0500

Msub= V*p*c= 50*1*0.1= 5 g

Where Msub- mass of (NH4)2SO4 , V - volume of solution, c - mass concentration of (NH4)2SO4